A loop of area 8.0cm is in constant magnetic field of B 0.15 T what is the magnetic flux through the loop when the loop is perendicular to the field?

1
Expert's answer
2021-12-20T12:36:19-0500

Magnetic flux is given as follows (https://en.wikipedia.org/wiki/Magnetic_flux):


"\Phi = BS\cos \theta"

where "B = 0.15T" is the mgnetic field, "S = 8cm^2 = 8\times 10^{-4}m^2" is the area of the loop, and "\theta =0" is the angle between loop's perpendicular and magnetic field. Thus, obtain:


"\Phi = 0.15\cdot 8\times 10^{-4} = 1.2\times 10^{-4}Wb"

Answer. "1.2\times 10^{-4}Wb".